OpenStudy (anonymous):

A snake starts moving at a speed of 2 m per second towards a target which is 5.8m away. In each second, the snake only moves two-thirds of the distance it traveled in the previous second. (a) Find the total distance that the snake has moves in the first five seconds. [Done; 5.21m] (b) Find the total distance that the snake can travel. [Done; 6m] >>>(c) When will the snake reach the target? <<<

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

the sum of distance, total distance = 5.8 so, plug in S= 5.8 in the equations you used, and keep time 't' as t only

hartnn (hartnn):

use the same formula as in part a)

OpenStudy (anonymous):

in part a, i work like this: 2+2*[(2/3)+(2/3)^2+...+(2/3)^4] =2+2*{[(2/3)(1-(2/3)^4]/[1-(2/3)]} =5.21

hartnn (hartnn):

so, for part c, you need to find number of terms, right ? let that be n 2+2* [(2/3)+(2/3)^2+....(2/3)^n] (2/3)+(2/3)^2+....(2/3)^n is the geometric series a1 = 2/3, r=2/3 so, instead of (2/3)+(2/3)^2+....(2/3)^n write \(\large a_1 \dfrac{1-r^n}{1-r}\)

OpenStudy (anonymous):

we need to find n. r=2/3. how about a1?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

a1=2m

OpenStudy (anonymous):

ohh yes

OpenStudy (anonymous):

but..

OpenStudy (anonymous):

whats the problem?

OpenStudy (anonymous):

(2/3)^n = 3/25 ?

OpenStudy (anonymous):

then use log

OpenStudy (anonymous):

...wait

OpenStudy (anonymous):

5.23??

OpenStudy (anonymous):

5.8 = {2(1-(2/3)^n)}/(1-(2/3))

OpenStudy (anonymous):

did u solve for n?

OpenStudy (anonymous):

n=5.23??

OpenStudy (anonymous):

\[2(\frac{ 1-(\frac{ 2 }{ 3 } )^{n}}{ 1-(\frac{ 2 }{ 3 }) })=5.28\]

OpenStudy (anonymous):

no...... it is 5.8

OpenStudy (anonymous):

yaya, typing mistake

OpenStudy (anonymous):

u used 5.28

OpenStudy (anonymous):

I got 8.39

OpenStudy (anonymous):

8.388

OpenStudy (anonymous):

got it now, thanks

OpenStudy (anonymous):

welcome